OpenStudy (anonymous):

solve for the variable. reduce the answer to lowest terms 5/6m=-30

OpenStudy (anonymous):

(6/5)*(5/6)m = -30*(6/5) m = -180/5

OpenStudy (anonymous):

m = -36